OpenStudy (anonymous):

why is the graph of y=log(10000x) the same as y=logx+4

OpenStudy (anonymous):

sorry i meant the graph of y=log(10000x) and y=logx+4

OpenStudy (sleepyjess):

\(\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\\\huge\cal\color{red}W\color{orange}E\color{goldenrod}L\color{yellow}C\color{lightgreen}O\color{darkgreen}M\color{turquoise}E~~\color{royalblue}T\color{purple}O~~\color{#00bfff}{Open}\color{#11c520}{Study}\\\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\)

OpenStudy (anonymous):

y = log(10000x) = log(10^4*x) = log(10^4)+log(x) = 4log(10)+log(x) = 4(1)+log(x) = 4+log(x) =log(x)+4

OpenStudy (anonymous):

They describe the same relationship between x and y

OpenStudy (anonymous):

damn son well in thanks a lot i knew they were the same just curious how mathematically jeeez perfect answer